solve for the square root of x+31=x+1

Answer by mananth(16949)

sqrt(x+31)=x+1
x+31= (x+1)^2
x+31= x^2+2x+1
x^2+x-30=0
x^2+6x-5x-30=0
x(x+6)-5(x+6)=0
(x+6)(x-5)=0
x=-6 or 5